# Further results for classical and universal characters twisted by roots of unity

### v1

@@ 4.3. $\mathrm{O}_{2tn+2}$ and $\operatorname{Sp}_{2tn}$
(4.5) $\mathrm{sp}_{\mu}(X,\omega X,\dots,\omega^{t-1}X)=(-1)^{\epsilon}\operatorname{sgn}(\sigma_{\mu})\,\mathrm{sp}_{\mu^{(t-1)}}(X^{t})\prod_{i=0}^{{(t-3)}/{2}}s_{(\mu^{(i)},-\mu^{(t-2-i)})_{2n}}(X^{t},{\overline{X}}^{t}),$
@@
$\epsilon=\displaystyle-\sum_{i=\left\lfloor{t}/{2}\right\rfloor}^{t-2}\binom{n_{i}(\mu,tn)+1}{2}+\frac{n(n+1)}{2}+nr,$
@@
(4.6) $\mathrm{o}^{\text{even}}_{\mu}(X,\omega X,\dots,\omega^{t-1}X,1)=(-1)^{\epsilon^{\prime}}\operatorname{sgn}(\sigma_{\mu})\,\mathrm{o}^{\text{even}}_{\mu^{(t-1)}}(X^{t},1)\prod_{i=0}^{{(t-3)}/{2}}s_{(\mu^{(i)},-\mu^{(t-2-i)})_{2n}}(X^{t},{\overline{X}}^{t}),$
@@
$\epsilon^{\prime}=\displaystyle\sum_{i=\left\lfloor{t}/{2}\right\rfloor}^{t-2}\binom{n_{i}(\mu,tn)}{2}.$
@@
The first factorization here uses Theorem 2.5. The second comes from Theorem 2.7, where we consider $\mu\in\mathcal{P}_{tn}$. Note that $n_{i}(\mu,tn)=n_{i+1}(\mu,tn+1)$, $20\leq i\leq t-$ 2, $n_{t-1}(\mu,tn)=n_{0}(\mu,tn+1)-1$. ∎
The results in this section suggest that a further generalization of [PW20, Conjecture 4] might be false.
@@ 5. Independence of twisted characters of variables
(5.1) $s_{\lambda}(X,Y)=\sum_{\mu\subset\lambda}s_{\lambda}(X)s_{\lambda/\mu}(Y).$
@@
Lemma 5.1 ([Kum22, Corollary 3.3]).
Let $\lambda$ and $\mu$ be partitions in $\mathcal{P}_{tm}$. Then the skew Schur polynomial $s_{\lambda/\mu}(Y,\omega Y,\dots,\omega^{t-1}Y)$ is given as follows.
@@ 5.1. Equality of hook Schur and Schur polynomials
In a separate vein, we have an interesting equality for skew Schur polynomials. We say that $\delta=(n,n-1,\dots,1)$ is a staircase partition.
@@
For $n\geq 1$, and all $\mu\subset\delta$, $s_{\delta/\mu}({X_{\infty}})=s_{\delta/\mu^{\prime}}({X_{\infty}}).$
@@
(5.10) $s_{(n+m,n+m-1,\dots,1)}(X,Y)=s_{(n,n-1,\dots,1)}(X)s_{(m,m-1,\dots,1)}(Y)\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j}),$
@@
(5.11) $s_{(n+m-1,n+m-2,\dots,1)}(X,Y)=s_{(n-1,n-2,\dots,1)}(X)s_{(m-1,m-2,\dots,1)}(Y)\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j}).$
@@ 6. Universal characters at roots of unity
$\mathfrak{so}_{\lambda^{(t/2-1)}}(1)=h_{\lambda^{(t/2-1)}}+h_{\lambda^{(t/2-1)}-1}=\begin{cases}1&\ell(\lambda^{(t/2-1)})=0,\\ 2&\ell(\lambda^{(t/2-1)})>0.\end{cases}$
@@
Suppose $\lambda$ is a partition in $\mathcal{P}_{t}$. Then $\mathfrak{o}_{\lambda}(1,\omega,\dots,\omega^{t-1})$ is nonzero if and only if $\mathrm{core}_{t}{(\lambda)}$ is a orthogonal $t$-core, $\ell(\lambda^{(i)})+\ell(\lambda^{(t-i)})\leq 1$ for all $1\leq i\leq\left\lfloor{t}/{2}\right\rfloor$ and $\lambda^{(0)}\in\{\emptyset,(1)\}$. In that case, if $\mathrm{core}_{t}{(\lambda)}$ has rank $r$,
$\mathfrak{o}_{\lambda}(1,\omega,\dots,\omega^{t-1})=(-1)^{\epsilon}\operatorname{sgn}(\sigma_{\lambda})\times\begin{cases}0&t\text{ even and }\lambda^{(t/2)}\neq\emptyset,\\ 1&t\text{ even and }\lambda^{(t/2)}=\emptyset,\\ 1&t\text{ odd},\end{cases}$
@@
$\mathfrak{o}_{\lambda}(1,\omega,\dots,\omega^{t-1})=(-1)^{\epsilon}\operatorname{sgn}(\sigma_{\lambda})\times\mathfrak{o}_{\lambda^{(0)}}(1)\times\prod_{i=1}^{\left\lfloor{(t-1)}/{2}\right\rfloor}\operatorname{rs}_{\lambda^{(i)},\lambda^{(t-i)}}(1)\times\begin{cases}\mathfrak{so}^{-}_{\lambda^{(t/2)}}(1)&t\text{ even},\\ 1&t\text{ odd},\end{cases}$
@@
$\mathfrak{so}_{\lambda^{((t-1)/2)}}(1)=h_{\lambda^{((t-1)/2)}}+h_{\lambda^{((t-1)/2)}-1}=\begin{cases}1&\ell(\lambda^{((t-1)/2)})=0,\\ 2&\ell(\lambda^{((t-1)/2)})>0.\end{cases}$
@@ 7. Classical characters at roots of unity
$\mathrm{sp}_{\lambda}(1,\omega,\dots,\omega^{t-1})=(-1)^{\epsilon}\operatorname{sgn}(\sigma_{\lambda})\,s_{\lambda^{(t-1)}}(1,1)\prod_{i=0}^{\left\lfloor{(t-3)}/{2}\right\rfloor}s_{\mu_{i}}(1,1)\times\begin{cases}s_{2\lambda^{(t/2-1)}}(1,1)&t\text{ even},\\ 1&t\text{ odd}.\end{cases}$
@@
$\mathrm{o}^{\text{even}}_{\lambda}(1,\omega,\dots,\omega^{t-1})=(-1)^{\epsilon}\operatorname{sgn}(\sigma_{\lambda})\,\mathrm{o}^{\text{even}}_{\lambda^{(0)}}(1)\prod_{i=1}^{\left\lfloor{(t-1)}/{2}\right\rfloor}s_{(\lambda^{(i)},-\lambda^{(t-i)})_{2}}(1,1)\\ \times\begin{cases}(-1)^{\lambda_{1}^{(t/2)}}\displaystyle{\operatorname{so}_{\lambda^{(t/2)}}(-1)}&t\text{ even},\\ 1&t\text{ odd},\end{cases}$
@@
$\mathrm{o}^{\text{even}}_{\lambda}(1,\omega,\dots,\omega^{t-1})=(-1)^{\epsilon}\operatorname{sgn}(\sigma_{\lambda})\,\prod_{i=1}^{\left\lfloor{(t-1)}/{2}\right\rfloor}s_{\mu_{i}}(1,1)\times\begin{cases}2&\lambda^{(0)}\neq\emptyset,\\ 1&\lambda^{(0)}=\emptyset.\end{cases}$
@@
$\operatorname{so}_{\lambda}(1,\omega,\dots,\omega^{t-1})=(-1)^{\epsilon}\operatorname{sgn}(\sigma_{\lambda})\,\prod_{i=0}^{\left\lfloor{(t-2)}/{2}\right\rfloor}s_{\mu_{i}}(1,1)\times\begin{cases}s_{2\lambda^{((t-1)/2)}}(1,1)&t\text{ odd},\\ 1&t\text{ even}.\end{cases}$
@@ Acknowledgements
We thank Dipendra Prasad for encouragement as well as many useful conversations, and Seamus Albion for discussions. The authors were partially supported by SERB Core grant CRG/2021/001592 as well as the DST FIST program - 2021 [TPN - 700661].

### v2

@@ 4.3. $\mathrm{O}({2tn+2})$ and $\operatorname{Sp}({2tn})$
(4.5) $\begin{split}\mathrm{sp}_{\mu}(X,\omega X,\dots,\omega^{t-1}X)=(-1)^{\epsilon}\operatorname{sgn}(\sigma_{\mu})\,\mathrm{sp}_{\mu^{(t-1)}}(X^{t})\prod_{i=0}^{{(t-3)}/{2}}&s_{(\mu^{(i)},-\mu^{(t-2-i)})_{2n}}(X^{t},{\overline{X}}^{t})\\ \times&{\begin{cases}\operatorname{so}_{\mu^{((t-2)/2)}}(X^{t})&t\text{ even},\\ 1&t\text{ odd,}\end{cases}}\end{split}$
@@
$\epsilon=\displaystyle-\sum_{i=\left\lfloor{t}/{2}\right\rfloor}^{t-2}\binom{n_{i}(\mu,tn)+1}{2}+\begin{cases}\frac{n(n+1)}{2}+nr&t\text{ even},\\ 0&t\text{ odd,}\end{cases}$
@@
[…]
@@
$\epsilon^{\prime}=\displaystyle\sum_{i=\left\lfloor{t}/{2}\right\rfloor}^{t-2}\Bigg(\binom{n_{i}(\mu,tn)}{2}+(t-1)n(n_{i}(\mu,tn)-n)\Bigg).$
@@
The first factorization here uses Theorem 2.5. The second comes from Theorem 2.7, where we consider $\mu\in\mathcal{P}_{tn}$. Note that $n_{i}(\mu,tn)=n_{i+1}(\mu,tn+1)$, $0\leq i\leq t-2$, $n_{t-1}(\mu,tn)=n_{0}(\mu,tn+1)-1$. ∎
The results in this section suggest that a further generalization of [PW22, Conjecture 4] might be false.
@@ 5. Independence of twisted characters of variables
(5.1) $s_{\lambda}(X,Y)=\sum_{\mu\subset\lambda}s_{{\mu}}(X)s_{\lambda/\mu}(Y).$
@@
Lemma 5.1.
Let $\lambda$ and $\mu$ be partitions in $\mathcal{P}_{tm}$. Then the specialized skew Schur polynomial $s_{\lambda/\mu}(Y,\omega Y,\dots,\omega^{t-1}Y)$ is given as follows.
@@
Proof.
By the Jacobi-Trudi type identity (2.6), we see that the required skew Schur polynomial is
(5.4) $s_{\lambda/\mu}(Y,\omega Y,\dots,\omega^{t-1}Y)=\det(h_{\beta_{i}(\lambda)-\beta_{j}(\mu)}(Y,\omega Y,\dots,\omega^{t-1}Y)).$
Permuting the rows and columns of the determinant by $\sigma_{\lambda}$ and $\sigma_{\mu}$ respectively, defined in (2.1), we see that the skew Schur polynomial is
$\operatorname{sgn}(\sigma_{\lambda})\operatorname{sgn}(\sigma_{\mu})\det\left(h_{\beta_{\sigma_{\lambda}(i)}(\lambda)-\beta_{\sigma_{\mu}(j)}(\mu)}(Y,\omega Y,\dots,\omega^{t-1}Y)\right),$
By Theorem 2.3 for $\lambda=(j)$, we have $h_{j}(Y,\omega Y,\dots,\omega^{t-1}Y)=0$ if $j\not\equiv 0\pmod{t}$ and $j>0$. Also, note that $h_{j}(Y,\omega Y,\dots,\omega^{t-1}Y)=0$ if $j<0$. Substituting these values in the above determinant, we see that the skew Schur polynomial is
$\operatorname{sgn}(\sigma_{\lambda})\operatorname{sgn}(\sigma_{\mu})\det\left(\begin{array}[]{ccccc}\widehat{S}_{0}&&&\\ &\widehat{S}_{1}&&\text{\huge 0}\\ &&\ddots\\ \text{\huge 0}&&&\widehat{S}_{t-1}\end{array}\right),$
where
$\widehat{S}_{p}=\left(h_{\beta^{(p)}_{i}(\lambda)-\beta^{(p)}_{j}(\mu)}(Y,\omega Y,\dots,\omega^{t-1}Y)\right)_{\begin{subarray}{c}1\leq i\leq n_{p}(\lambda)\\ 1\leq j\leq n_{p}(\mu)\end{subarray}},\quad p\in[0,t-1].$
If $\mathrm{core}_{t}{(\lambda)}\neq\mathrm{core}_{t}{(\mu)}$, then $n_{i}(\lambda)\neq n_{i}(\mu)$ for some $i\in[0,t-1]$ and the $(i+1)$’th diagonal block is not a square block. So, $s_{\lambda/\mu}(Y,\omega Y,\dots,\omega^{t-1}Y)=0.$ If $\mathrm{core}_{t}{(\lambda)}=\mathrm{core}_{t}{(\mu)}$, then $n_{i}(\lambda)=n_{i}(\mu)$ for all $0\leq i\leq t-1$. In this case, $\det(\widehat{S}_{p})=s_{\lambda^{(p)}/\mu^{(p)}}(Y^{t})$. This gives the desired result. ∎
@@ 5.1. Equality of hook Schur and Schur polynomials
In a separate vein, we have an interesting equality for skew Schur polynomials. We say that ${\delta_{n}}=(n,n-1,\dots,1)$ is a staircase partition.
@@
For $n\geq 1$, and all $\mu\subset{\delta_{n}}$, $s_{{\delta_{n}}/\mu}({X_{\infty}})=s_{{\delta_{n}}/\mu^{\prime}}({X_{\infty}}).$
@@
(5.11) $s_{{\delta_{n+m}}}(X,Y)=s_{{\delta_{n}}}(X)s_{{\delta_{m}}}(Y)\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j}),$
@@
(5.12) $s_{{\delta_{n+m-1}}}(X,Y)=s_{{\delta_{n-1}}}(X)s_{{\delta_{m-1}}}(Y)\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j}).$
@@ 6. Universal characters at roots of unity
$\mathfrak{so}_{\lambda^{(t/2-1)}}(1)={h_{\lambda^{(t/2-1)}}(1)+h_{\lambda^{(t/2-1)}-1}(1)}=\begin{cases}1&\ell(\lambda^{(t/2-1)})=0,\\ 2&\ell(\lambda^{(t/2-1)})>0.\end{cases}$
@@
Suppose $\lambda$ is a partition in $\mathcal{P}_{t}$. Then $\mathfrak{o}_{\lambda}(1,\omega,\dots,\omega^{t-1})$ is nonzero if and only if $\mathrm{core}_{t}{(\lambda)}$ is an orthogonal $t$-core, $\ell(\lambda^{(i)})+\ell(\lambda^{(t-i)})\leq 1$ for all $1\leq i\leq\left\lfloor{t}/{2}\right\rfloor$ and $\lambda^{(0)}\in\{\emptyset,(1)\}$. In that case, if $\mathrm{core}_{t}{(\lambda)}$ has rank $r$,
[…]
@@
[…]
@@
$\mathfrak{so}_{\lambda^{((t-1)/2)}}(1)={h_{\lambda^{((t-1)/2)}}(1)+h_{\lambda^{((t-1)/2)}-1}(1)}=\begin{cases}1&\ell(\lambda^{((t-1)/2)})=0,\\ 2&\ell(\lambda^{((t-1)/2)})>0.\end{cases}$
@@ 7. Classical characters at roots of unity
$\begin{split}\mathrm{sp}_{\lambda}(1,\omega,\dots,\omega^{t-1})=(-1)^{\epsilon}\operatorname{sgn}(\sigma_{\lambda})\,s_{\lambda^{(t-1)}}(1,1)\prod_{i=0}^{\left\lfloor{(t-3)}/{2}\right\rfloor}&s_{{(\lambda^{(i)},-\lambda^{(t-2-i)})_{2}}}(1,1)\\ &\times\begin{cases}s_{2\lambda^{(t/2-1)}}(1,1)&t\text{ even},\\ 1&t\text{ odd}.\end{cases}\end{split}$
@@
[…]
@@
$\mathrm{o}^{\text{even}}_{\lambda}(1,\omega,\dots,\omega^{t-1})=(-1)^{\epsilon}\operatorname{sgn}(\sigma_{\lambda})\,\prod_{i=1}^{\left\lfloor{(t-1)}/{2}\right\rfloor}s_{{(\lambda^{(i)},-\lambda^{(t-i)})_{2}}}(1,1)\times\begin{cases}2&\lambda^{(0)}\neq\emptyset,\\ 1&\lambda^{(0)}=\emptyset.\end{cases}$
@@
$\operatorname{so}_{\lambda}(1,\omega,\dots,\omega^{t-1})=(-1)^{\epsilon}\operatorname{sgn}(\sigma_{\lambda})\,\prod_{i=0}^{\left\lfloor{(t-2)}/{2}\right\rfloor}s_{{(\lambda^{(i)},-\lambda^{(t-1-i)})_{2}}}(1,1)\times\begin{cases}s_{2\lambda^{((t-1)/2)}}(1,1)&t\text{ odd},\\ 1&t\text{ even}.\end{cases}$
@@ Acknowledgements
We thank Dipendra Prasad for encouragement as well as many useful conversations, and Seamus Albion for discussions. We also thank the referees for their many useful comments. The authors were partially supported by SERB Core grant CRG/2021/001592 as well as the DST FIST program - 2021 [TPN - 700661].